# Diverging Towards Hallucination: Detection of Failures in Vision-Language Models via Multi-token Aggregation

### v1

@@ Abstract
Vision–language models (VLMs) now rival human performance on many multimodal tasks, yet they still hallucinate objects or generate unsafe text. Current hallucination detectors, e.g., single-token linear probing (SLP) and $P(\text{True})$, typically analyze only the logit of the first generated token—or just its highest-scoring component—overlooking richer signals embedded within earlier token distributions. We demonstrate that analyzing the complete sequence of early logits potentially provides substantially more diagnostic information. We emphasize that hallucinations may only emerge after several tokens, as subtle inconsistencies accumulate over time. By analyzing the Kullback–Leibler (KL) divergence between logits corresponding to hallucinated and non-hallucinated tokens, we underscore the importance of incorporating later-token logits to more accurately capture the reliability dynamics of VLMs. In response, we introduce Multi-Token Reliability Estimation (MTRE), a lightweight, white-box method that aggregates logits from the first ten tokens using multi-token log-likelihood ratios and self-attention. Despite the challenges posed by large vocabulary sizes and long logit sequences, MTRE remains efficient and tractable. On MAD-Bench, MM-SafetyBench, MathVista, and four compositional-geometry benchmarks, MTRE improves AUROC by 9.4 ± 1.3 points over SLP and by 12.1 ± 1.7 points over $P(\text{True})$, setting a new state-of-the-art in hallucination detection for open-source VLMs.
@@ 1 Introduction
Vision-language models (VLMs) have recently achieved groundbreaking performance across a range of multimodal tasks, from image captioning to visual question answering. Despite these advances, VLMs remain susceptible to generating hallucinated, unsafe, or contextually inappropriate outputs, particularly when faced with ambiguous or adversarial inputs. Such vulnerabilities pose serious challenges for deploying these models in real-world, safety-critical applications. For deep-learning in general, significant research efforts have been devoted to improving model calibration and quantifying uncertainty (Guo et al., 2017; Gal and Ghahramani, 2016; Kendall and Gal, 2017). However, many of these traditional approaches treat VLMs as black boxes, relying solely on output-level statistics without tapping into the rich internal representations that these models naturally generate.
@@
The current practice to address hallucination in VLMs relies directly on the logits associated with generated tokens Steyvers et al. (2025). Intuitively, this method assumes that higher model confidence in generating a token implies a lower likelihood of hallucination. More interestingly, a recent study by Zhao et al. (2024) demonstrated that the logit of the first token in an output sequence alone contains sufficient information to assess the reliability of the generated text. Our work challenges these viewpoints: we argue that focusing exclusively on the confidence or a single token inherently limits the contextual information available, resulting in suboptimal hallucination detection. In particular, we leverage the potential connection between KL divergence and class separation to highlight the importance of utilizing later-generated logits in the reliability of VLMs (Sect. 4). Our hypothesis is, once a hallucinated token is: produced, the corresponding generated logit and/or surrounding logits will consequently shift away from the the model’s prior belief of the environment, which directly translates to a higher divergence. However, as directly computing divergence from the model’s prior belief is prohibitive due to the requirement of the prior, we derive a relative measure and directly compare between hallucination and non-hallucination scenarios. Our empirical results confirm that the occurrence of a hallucination at a particular token position does lead to a noticeable divergence. Additionally, we observe that when this divergence emerges at later token positions, the effectiveness of hallucination detection based solely on the initial token logits Zhao et al. (2025) often significantly deteriorates compared to their performance when divergence occurs around earlier tokens. This finding suggests that later tokens may contain critical reliability-related information absent in earlier tokens. Consequently, we propose and develop a detection method (Sect. 5) leveraging logits from multiple output tokens, capturing a richer and more nuanced representation of the model’s internal decision-making process. Extensive experiments (Sect. 6) on benchmark datasets such as MAD-Bench Li et al. (2023), MM-SafetyBench Liu et al. (2023a), MathVista Lu et al. (2023), and other various arithmetic centered questions Rahmanzadehgervi et al. (2024) demonstrate that our approach allows for the usage of more tokens to potentially improve reliability prediction metrics, thereby establishing a practical and computationally efficient pathway for enhancing the safety of VLM outputs.
@@ 2 Related Work
The reliability of deep neural networks has been extensively studied through the lens of calibration and uncertainty quantification.Guo et al. (2017) provided a seminal analysis demonstrating that high-performing neural networks often exhibit poor calibration, and introduced temperature scaling as a simple yet effective post-hoc adjustment. Complementing these findings, Bayesian methods have been leveraged to estimate uncertainty in neural networks. Gal and Ghahramani (2016) pioneered the use of dropout as a Bayesian approximation, and Kendall and Gal (2017) further developed frameworks to jointly model aleatoric and epistemic uncertainty, especially within computer vision tasks.
@@
In parallel, the self-assessment capabilities of large language models (LLMs) have garnered significant attention. Recent studies demonstrate that prompting LLMs to output confidence scores (often quantified via the P(true) metric Kadavath et al. (2022a)) can provide a proxy for prediction reliability. However, these methods typically treat the model as a black box, focusing solely on output-level probabilities rather than the underlying internal representations.
@@
A related stream of research investigates semantic uncertainty using loss-based measures. For example, there have been efforts to utilize semantic loss metrics to capture the inherent ambiguity in model outputs Grewal et al. (2024). While these approaches yield important insights into output variability, they do not exploit the fine-grained, white-box information available during the early stages of sequence generation.
Retrieval-Augmented Generation (RAG) has shown significant promise in enhancing Large Language Models (LLMs) by enabling them to access and incorporate relevant external knowledge during text generation. This approach improves factual accuracy and mitigates the limitations of static model parameters Ayala and Bechard (2024). However, applying RAG in the context of Visual Question Answering (VQA) Antol et al. (2015) presents unique challenges. Unlike purely textual tasks, VQA requires the model to interpret visual inputs alongside natural language queries, making the integration of retrieved textual information more complex. The retrieved documents may not align well with the visual content, leading to difficulties in grounding the information effectively. Moreover, fusing multimodal data (image, question, and retrieved text) in a coherent manner remains an open research problem, limiting the direct applicability of RAG techniques in VQA scenarios.
@@
More recently, Zhao et al. (2025) demonstrated that the logit distribution of the very first token in VLM outputs encodes latent signals related to model behavior and reliability. This finding suggests that internal representations carry richer information than what is apparent from the final output alone. However, the focus on a single token may overlook additional contextual cues. In contrast, our approach aggregates embeddings from the first ten tokens, thereby capturing a more nuanced and comprehensive snapshot of the model’s internal state. Our work synthesizes and extends prior research in calibration, Bayesian uncertainty, and semantic uncertainty. By leveraging white-box access to early token embeddings, we provide a rigorous framework that not only enhances predictive performance but also deepens our understanding of the internal mechanisms governing VLM behavior.
3 Background
@@
To rigorously understand and detect hallucinations in VLMs, we first clarify the underlying autoregressive generation mechanism used by these models and introduce Kullback-Leibler Divergence as a tool for quantifying model uncertainty and unexpectedness. This background provides the essential framework needed to motivate and justify our multi-token reliability estimation method.
3.1 Autoregressive Generation and Early-Token Representations
@@
A VLM with parameters $\theta$ processes multimodal inputs, typically comprising an image $x\in\mathcal{X}$ and a text-based prompt represented as a token sequence $t=(t_{1},t_{2},\dots,t_{M})$, where each token $t_{i}\in\mathcal{V}$ and $\mathcal{V}$ is a finite vocabulary. Given these inputs, the VLM generates an output token sequence $y=(y_{1},y_{2},\dots,y_{K})$ autoregressively.
$P_{\theta}(y\mid x,t)\;=\;\prod_{k=1}^{K}P_{\theta}\!\bigl{(}y_{k}\mid x,t,y_{<k}\bigr{)},\qquad y_{<k}:=(y_{1},\dots,y_{k-1}).$ (1)
@@
At each generation step $k$, the model estimates the conditional probability distribution of the next token based on previously generated tokens and input context:
$P_{\theta}(y_{k}|x,t,y_{<k})=\text{softmax}(f_{\theta}(x,t,y_{<k}))$ (2)
Here, the function $f_{\theta}(x,t,y_{<k})$ produces logits $\ell_{k}\in\mathbb{R}^{|\mathcal{V}|}$, representing unnormalized probabilities over the vocabulary. Specifically, logits $\ell_{k}$ at generation step $k$ are given by:
$\ell_{k}=f_{\theta}(x,t,y_{<k})\quad\text{with}\quad\ell_{k}\in\mathbb{R}^{|\mathcal{V}|}$ (3)
Typically, the VLM employs sampling strategies (e.g., greedy decoding, beam search, or nucleus sampling) to select tokens from the computed probability distributions. Thus, the generated token at step $k$ is determined as:
$y_{k}=g(\ell_{k}),\quad\text{where}\quad g:\mathbb{R}^{|\mathcal{V}|}\rightarrow\mathcal{V}$ (4)
The first logit $\ell_{1}$ thus encodes the model’s initial alignment between the multimodal prompt and the language head. Empirically, linear probes on $\ell_{1}$ already reveal a substantial amount of reliability signal (Zhao et al., 2024; Kadavath et al., 2022b). However, hallucinations may emerge after the first token, once the model conditions on its own (possibly flawed) partial output.
3.2 Kullback-Leibler (KL) divergence
The Kullback-Leibler (KL) divergence is a measure of how one probability distribution differs from a second, reference probability distribution. Formally, for continuous probability distributions $P$ and $Q$ defined on the same probability space, the KL divergence from $Q$ to $P$ is defined as:
$D_{\text{KL}}(P\;||\;Q)=\int_{x}P(x)\,\log\frac{P(x)}{Q(x)}\,dx$
For discrete probability distributions, the integral is replaced with a sum:
$D_{\text{KL}}(P\;||\;Q)=\sum_{x}P(x)\,\log\frac{P(x)}{Q(x)}$
In information theory, KL divergence represents the expected excess surprise from using $Q$ as a model when the true distribution is $P$. It can be interpreted as the amount of information lost when $Q$ is used to approximate $P$. In this setting, we primarily utilize the discrete variation.
Figure 1: We show two types of divergence: (1) Divergence between responses with different labels for the final classification tasks discussed in Zhao et al. (2025) (solid lines), and (2) Divergence between hallucinated and non-hallucinated samples during the self-evaluation task discussed in Section 6 (dashed lines). We visualize divergence for each label set to allow for detailed analysis. Notably, divergence in self-evaluation tasks tends to peak much later than in final classification tasks. We hypothesize that the position and magnitude of the divergence peak relative to the first token may inform improvements to the first-token linear probing technique. Our results (Table 1) suggest that linear probing shows strong performance on tasks represented by solid lines.
4 Class Separation between Hallucinations and Non-Hallucinations in VLMs
Consider the task of classifying conditional distribution of the $i$-th logit under the non-hallucination and hallucination classes from a vision-language model (VLM), specifically the logits corresponding to:
$P_{i}=P(D^{\text{hallu}}_{i}\mid D_{<i},\text{prompt})\text{ and }Q_{i}=P(D^{\text{non-hallu}}_{i}\mid D_{<i},\text{prompt}),$
For the purposes of hallucination detection, we hypothesize that the Kullback–Leibler (KL) divergence between these two distributions is correlated to the degree of class separability in the context of this binary classification task, i.e:
$Corr(D_{KL},\phi)<0$
Where $D_{KL}=KL[P_{i}||Q_{i}],$ and $\phi$ is the model’s binary misclassification rate.
x In contrast to the tasks used by Zhao et al. (2025), we explore this hypothesis using tasks that allow hallucinations to emerge later in a model’s prediction as seen in Figure 2—potentially triggered by multimodal ambiguity, limited model robustness, or internal inconsistencies that compound during autoregressive generation Xu et al. (2025). For a method that is trained to classify if a sentence is hallucinated or not, we expect that the level of divergence between hallucinated and non hallucinated tokens can play a role in improving the models performance.
Figure 2: We show the divergence between hallucinated and non-hallucinated samples in the Arithmetic setting (Type 1 responses 4). We again visualize divergence from each label set to allow for detailed analysis.
Although computing the exact correlation to performance is challenging—due to differing performance scales, non-linear interactions, and the high-dimensional nature of model behavior—the pronounced divergence observed in later parts of the sentence highlights the need for a method capable of leveraging tokens from regions of high input divergence.
5 Method
Our empirical studies show that hallucinations may not always reveal themselves in the very first token of a model’s answer; instead they might emerge once progressive inconsistencies cumulate across successive tokens. Unfortunately, designing a method capable of making predictions over extended logit sequences—often associated with large vocabulary sizes ($\mathrm{R}^{32,000})$—is challenging, as it can quickly become intractable due to memory constraints.
Given hardware limitations and the necessity to make predictions over the entire sentence, we introduce Multi-Token Reliability Estimation (MTRE), a computationally efficient procedure capable of utilizing the entire trajectory of model outputs for detecting hallucinations in VLMs. MTRE re-casts hallucination detection as a sequential log-likelihood–ratio test Wald (1992) that accumulates evidence over the first $k$ generated logits (typically $k\!=\!10$ for our experiments).
5.1 Probabilistic Model
For a given sample let:
$\mathbf{X}=[x_{0},x_{1},\dots,x_{T-1}]\in\mathbb{R}^{T\times d}$

### v2

# MTRE: Multi-Token Reliability Estimation for Hallucination Detection in VLMs

@@ Abstract
Vision–language models (VLMs) now rival human performance on many multimodal tasks, yet they still hallucinate objects or generate unsafe text. Current hallucination detectors, e.g., single-token linear probing (LP) and $P(\text{True})$, typically analyze only the logit of the first generated token—or just its highest-scoring component—overlooking richer signals embedded within earlier token distributions. We demonstrate that analyzing the complete sequence of early logits potentially provides substantially more diagnostic information. We emphasize that hallucinations may only emerge after several tokens, as subtle inconsistencies accumulate over time. By analyzing the Kullback–Leibler (KL) divergence between logits corresponding to hallucinated and non-hallucinated tokens, we underscore the importance of incorporating later-token logits to more accurately capture the reliability dynamics of VLMs. In response, we introduce Multi-Token Reliability Estimation (MTRE), a lightweight, white-box method that aggregates logits from the first ten tokens using multi-token log-likelihood ratios and self-attention. Despite the challenges posed by large vocabulary sizes and long logit sequences, MTRE remains efficient and tractable. Across MAD-Bench, MM-SafetyBench, MathVista, and four compositional-geometry benchmarks, MTRE achieves a 9.4% gain in Accuracy and a 14.8% gain in AUROC over standard detection methods, establishing a new state of the art in hallucination detection for open-source VLMs.
@@ 1 Introduction
Vision-language models (VLMs) have recently achieved groundbreaking performance across a range of multimodal tasks, from image captioning to visual question answering. Despite these advances, VLMs remain susceptible to generating hallucinated, unsafe, or contextually inappropriate outputs, particularly when faced with ambiguous or adversarial inputs. Such vulnerabilities pose serious challenges for deploying these models in real-world, safety-critical applications. For deep-learning in general, significant research efforts have been devoted to improving model calibration and quantifying uncertainty (Guo et al., 2017; Gal & Ghahramani, 2016; Kendall & Gal, 2017). However, many of these traditional approaches treat VLMs as black boxes, relying solely on output-level statistics without tapping into the rich internal representations that these models naturally generate.
@@
The current practice to address hallucination in VLMs relies directly on the logits associated with generated tokens (Steyvers et al., 2025). Intuitively, this method assumes that higher model confidence in generating a token implies a lower likelihood of hallucination. More interestingly, a recent study by (Zhao et al., 2025) demonstrated that the logit of the first token in an output sequence alone contains sufficient information to assess the reliability of the generated text. Our work challenges these viewpoints: we argue that focusing exclusively on the confidence or a single token inherently limits the contextual information available, resulting in suboptimal hallucination detection. In particular, we leverage the potential connection between KL divergence and class separation to highlight the importance of utilizing later-generated logits in the reliability of VLMs (Sect. 3.2). Our hypothesis is, once a hallucinated token is: produced, the corresponding generated logit and/or surrounding logits will consequently shift away from the the model’s prior belief of the environment, which directly translates to a higher divergence. However, as directly computing divergence from the model’s prior belief is prohibitive due to the requirement of the prior, we derive a relative measure and directly compare between hallucination and non-hallucination scenarios. Our empirical results confirm that the occurrence of a hallucination at a particular token position does lead to a noticeable divergence. Additionally, we observe that when this divergence emerges at later token positions, the effectiveness of hallucination detection based solely on the initial token logits (Zhao et al., 2025) often significantly deteriorates compared to their performance when divergence occurs around earlier tokens. This finding suggests that later tokens may contain critical reliability-related information absent in earlier tokens. Consequently, we propose Multi-Token Reliability Estimation (MTRE) (Sect. 4), along with several variants, which leverage logits from multiple output tokens to capture a richer and more nuanced representation of the model’s internal decision-making process.
Figure 1: Summary of experiments on MAD-Bench and $M$-Safety (5 methods on 2 detection tasks on 4 VLMs in 2 datasets): Each cell shows the fraction of experiments where the method in the row outperforms the method in the column measured by Accuracy and AUROC, respectively.
Figure 1 highlights our key results, showing the significant performance gains achieved by the proposed MTRE method. Unlike approaches that rely solely on the first token, MTRE aggregates information across multiple tokens, leading to more robust predictions. Extensive experiments (Sect. 5) on benchmark datasets, including MAD-Bench (Li et al., 2023), MM-SafetyBench (Liu et al., 2023b), MathVista (Lu et al., 2023), and a variety of arithmetic-focused questions (Rahmanzadehgervi et al., 2024), demonstrate that leveraging multiple tokens leads to more reliable hallucination detection. This establishes a practical and computationally efficient pathway for enhancing the safety of VLM.
@@ 2 Related Work
The self-assessment capabilities of VLMs have garnered significant attention with many preliminary techniques have come about. Several strategies have been proposed to address the challenges of mitigating or detecting hallucination. One direct approach is to align VLMs with human preferences through Reinforcement Learning with Human Feedback (Chen et al., 2023). Another is to curate datasets containing both harmful and benign samples and finetune an LLM to detect unsafe content (Pi et al., 2024). However, both approaches demand substantial computational resources and have shown potential of inducing catastrophic forgetting (Mukhoti et al., 2024). Prompt tuning (Yao et al., 2023), either through manual design or automated learning of task-specific prompts, has also been explored. While useful, this method tends to be suboptimal: manual design is non-trivial, and automated prompt learning for VLMs is computationally expensive. There have been some initial works that utilize the image directly, such as (Kiana Avestimehr & Mushtaq, 2025), which aims to estimate visual uncertainty by leveraging visual contrast between an observation with task-relevant features and one without; however, this requires knowledge of the task-relevant features, which is not always available. In addition, other common uncertainty quantification techniques for VLMs (Kostumov et al., 2024) require reformatting the prompt in the form of a multiple choice question, which, unfortunately, for open-ended responses, may alter the true uncertainty of the original context (Kumar et al., 2023). Another line of work leverages auxiliary models to guide uncertainty estimation (Duan et al., 2024), but this introduces an external dependency that may limit scalability and robustness. Sampling-based approaches (Orgad et al., 2025; Kuhn et al., 2023) have also been investigated, but inference constraints may be restrictive to one sample, and methods may be sensitive to sampling variance. Moreover, many previously successful auto-regressive uncertainty methods (Malinin & Gales, 2021) have not yet demonstrated scalability to the large models used today.
@@
Recent studies demonstrate that prompting LLMs to output confidence scores (often quantified via the P(true) uncertainty score (Steyvers et al., 2025; Kadavath et al., 2022a)) can provide a proxy for prediction reliability. However, these methods typically treat the model as a black box, focusing solely on output-level probabilities rather than the underlying internal representations.
@@
A related stream of research investigates semantic uncertainty using loss-based measures. For example, there have been efforts to utilize semantic loss metrics to capture the inherent ambiguity in model outputs (Grewal et al., 2024). While these approaches yield important insights into output variability, they do not exploit the fine-grained, white-box information available during the early stages of sequence generation.
@@
More recently, Zhao et al. (2025) demonstrated that the logit distribution of the very first token in VLM outputs encodes latent signals related to model behavior and reliability. This finding suggests that internal representations carry richer information of the image and text than what is apparent from the final output alone. However, the focus on a single token may overlook additional contextual cues. In contrast, our approach aggregates embeddings from the first N tokens, thereby capturing a more nuanced and comprehensive snapshot of the model’s internal state. Our work synthesizes and extends prior research in calibration, Bayesian uncertainty, and semantic uncertainty. By leveraging white-box access to early token embeddings, we provide a rigorous framework that not only enhances predictive performance but also deepens our understanding of the internal mechanisms governing VLM behavior.
3 Preliminaries
@@
To investigate and detect hallucinations using logits, we first clarify the autoregressive generation mechanism underlying VLMS. We then introduce Kullback–Leibler Divergence as a tool for quantifying differences in model behavior between hallucinated and non-hallucinated generations. These preliminary insights not only provide motivation but also guide the design of our multi-token reliability estimation method.
3.1 Autoregressive Generation in VLMs
@@
A VLM $f$ with parameters $\theta$ processes multimodal inputs, typically comprising an image $x\in\mathcal{X}$ and a text-based prompt represented as a token sequence $\delta=(\delta_{1},\delta_{2},\dots,\delta_{M})$ , where each token $\delta_{i}\in\mathcal{V}$ and $\mathcal{V}$ is a finite vocabulary. Given these inputs, the VLM generates an output token sequence $y=(y_{1},y_{2},\dots,y_{T})$ autoregressively by estimating the joint probability:
$P(y\mid x,\delta)\;=\;\prod_{t=1}^{T}P\!\bigl(y_{t}\mid x,\delta,y_{<t}\bigr),\qquad y_{<t}:=(y_{1},\dots,y_{t-1}).$ (1)
@@
Specifically, at each generation step $k$, the model estimates the conditional probability distribution of the next token based on previously generated tokens and input context:
$P(y_{t}|x,\delta,y_{<t})\approx\mathrm{softmax}(f_{\theta}(x,\delta,y_{<t})).$ (2)
The VLM’s output $\ell_{t}=f_{\theta}(x,t,y_{<t})\in\mathbb{R}^{|\mathcal{V}|}$ is called logits, representing unnormalized probabilities over the vocabulary. Given $\{\ell_{t}\}_{t=1}^{T}$, sampling strategies (e.g., greedy decoding, beam search, or nucleus sampling) are employed to produce tokens from the computed probability.
3.2 Token-wise Divergence between Hallucinations and Non-Hallucinations
Several recent works (Zhao et al., 2025; Kadavath et al., 2022b) suggest that the first logit, $\ell_{1}$, encodes the model’s initial alignment between the multimodal prompt and the language head. Empirically, linear probing on $\ell_{1}$ Gurnee & Tegmark (2023) has been shown to perform well for hallucination detection. This observation is consistent with findings that the distribution of the very first token is particularly informative for predicting model behavior. However, since the model conditions on its own (potentially flawed) generations, hallucinations can potentially emerge after the first step, and subsequent logits tend to be less discriminative than the first logit.
Figure 2: We measure the KL divergence between the conditional probability distributions of the next token under hallucinated versus non-hallucinated generations, i.e., $\mathrm{softmax}(\ell_{t})$ when $y_{t}$ is hallucinated versus $\mathrm{softmax}(\ell_{t})$ when $y_{t}$ is non-hallucinated, in the Type 1 classification tasks and Type 2 self-evaluation tasks among different models and datasets.
To validate this hypothesis, we compare the estimated probability of a token when it is hallucinated versus when it is not, i.e., $P^{\text{hallu}}_{t}:=P(y_{t}\mid x,\delta,y_{<t},\,y_{t}\text{ is hallucinated})$ versus $P^{\text{non-hallu}}_{t}:=P(y_{t}\mid x,\delta,y_{<t},\,y_{t}\text{ is non-hallucinated})$. In particular, we measure the KL divergence:
$\mathcal{D}_{t}:=D_{\mathrm{KL}}\!\left(P^{\text{hallu}}_{t}\|P^{\text{non-hallu}}_{t}\right)\;=\;\sum_{v\in\mathcal{V}}P^{\text{hallu}}_{t}(v)\,\log\frac{P^{\text{hallu}}_{t}(v)}{P^{\text{non-hallu}}_{t}(v)}.$
Figure 3: The KL divergence between hallucinated and non-hallucinated responses in the Arithmetic dataset (Type 1).
In Figures 2 and 3, we compute the KL divergence $\mathcal{D}_{t}$ at different positions $t$ for hallucinated versus non-hallucinated responses across two VLMs, evaluated on the Safety (Liu et al., 2023b), MAD (Li et al., 2023), and Arithmetic Rahmanzadehgervi et al. (2024) benchmarks (Details of the experiments are provided in C.1). Unlike Zhao et al. (2025), which focuses primarily on hallucinations in direct model outputs (Type 1), our analysis also considers self-evaluation tasks (Type 2), where hallucinations typically arise later in the response. As shown in Figure 2, divergences in Type 2 tasks tend to emerge at later token positions compared to Type 1 classification tasks. A similar divergence pattern in KL divergence is also observed for Type 1 hallucinations in Arithmetic tasks (Figure 3), where critical information often occurs toward the end of the model’s response. Intuitively, a sharp increase in KL divergence at a given token position indicates the onset of hallucination. Consequently, relying on logits from earlier positions is likely to be suboptimal for detection.
In fact, the Token-wise divergence behavior observed in Figures 2 and 3 suggests the limitations of relying solely on $\ell_{1}$ for hallucination detection. The observation strongly supports and explains why integrating more tokens can lead to significant detection gain, as will be shown in Section 5.3. Especially, experimental results in Figure 5 and Table 2 confirm that probing methods restricted to the first token (Zhao et al., 2025) are suboptimal for Type 2 settings, where hallucinations appear later in the sequence. A similar low detection performance of depending only on the first token (Table 1) is also observed for Type 1 hallucinations in Arithmetic tasks, where critical information often occurs toward the end of the model’s response. Collectively, these results highlight the need for more comprehensive detection strategies that incorporate information from multiple tokens to improve robustness.
4 Multi-Token Reliability Estimation
Algorithm 1 Sentence-Level evidence aggregation for MTRE
1:Test subset $\mathbb{S}_{\text{test}}$, trained $f_{\theta}$
2:Sentence predictions $\hat{Y}_{s_{i}}$
3:for each sentence $s_{i}\in\mathbb{S}_{\text{test}}$ do
4:  for $t=1\to\tau_{i}\leq T_{i}$ do
5:   $p^{s_{i}}_{t}\leftarrow f_{\theta}(x^{s_{i}}_{t})$
6:   $z_{s_{i},t}\leftarrow\log\tfrac{p^{s_{i}}_{t}}{1-p^{s_{i}}_{t}}$
7:   $L_{s_{i},\tau_{i}}\leftarrow L_{s_{i},\tau_{i}}+z_{s_{i},t}$
8:  end for
9:  $\hat{Y}_{s_{i}}\leftarrow\begin{cases}1&L_{s_{i},\tau_{i}}\geq 0\\ 0&\text{otherwise}\end{cases}$
10:end for
Algorithm 2 LLR collection for MTRE
1:Fold subset $\text{fold}_{j}$, trained $f_{\theta}$, optional initial LLR Pair Dataset from past folds $\mathcal{D}_{0}$
2:LLR Pair Dataset $\mathcal{D}$
[…]
4:for each sentence $s_{i}\in\text{fold}_{j}$ do
5:  for $t=1\to\tau_{i}\leq T_{i}$ do
6:   $p^{s_{i}}_{t}\leftarrow f_{\theta}(x^{s_{i}}_{t})$
7:   $z_{i,t}\leftarrow\log\tfrac{p^{s_{i}}_{t}}{1-p^{s_{i}}_{t}}$
8:   Add $(z_{i,t},Y_{s_{i}})$ to $\mathcal{D}$
9:  end for
10:end for
11:return $\mathcal{D}$
Hallucinations in VLMs often emerge progressively: early tokens may look plausible while inconsistencies accrete across subsequent tokens (demonstrated in Figure 2 and 3). Detecting such failures, therefore, benefits from multi-token evidence rather than single-token probes. However, scoring long sequences with large vocabularies can be memory- and latency-limited. We address this by using a short prefix (typically the first T=10T{=}10 tokens) and by designing a Multi-Token Reliability Estimation (MTRE) procedure that is both statistically principled and computationally light. At a high level, MTRE trains and applies a reliability classifier $f_{\theta}$ to detect hallucination signatures at the token level. This design is motivated by the strong performance of first-token methods in certain tasks, suggesting that logits’ internal values contain rich signals for hallucination detection. Building on this insight, MTRE formulates hallucination detection as a calibrated sequential log-likelihood–ratio (LLR) test (Wald, 1992), incorporating (i) token-level aggregation, (ii) adaptive early stopping, and (iii) out-of-fold calibration.
4.1 Token Level Training for Reliability Classifier
For a given sentence $s_{i}\in\mathbb{S}$, let
$\mathbf{X}_{s_{i}}=[x^{s_{i}}_{0},x^{s_{i}}_{1},\dots,x^{s_{i}}_{T_{i}}]\in\mathbb{R}^{T_{i}\times d}$